OpenStudy (azureilai):

If a snowball melts so that its surface area decreases at a rate of 1 cm^2/min, find the rate at which the diameter decreases when the diameter is 10 cm. I'm not sure where to start with this one. Help please. I'm using "S" to represent surface area and "x" for diameter.

OpenStudy (baru):

whats the formula for surface are of a sphere?

OpenStudy (baru):

*surface area

OpenStudy (azureilai):

4*pi*r^2

OpenStudy (baru):

can you re-write the formula in terms of diameter?

OpenStudy (azureilai):

4*pi*(d/2)^2

OpenStudy (azureilai):

4*pi*(x/2)^2 (I forgot I was using x to represent diameter)

OpenStudy (baru):

ok \[s=\pi (x/2)^2\\simplify\\s=\pi x^2\]

OpenStudy (baru):

okay?

OpenStudy (azureilai):

Where did the 4 go?

OpenStudy (baru):

\[s=\frac{4\pi x^2}{4}\] i have just expanded the formula you gave

OpenStudy (baru):

4 and 4 cancel

OpenStudy (azureilai):

ok. I see.

OpenStudy (baru):

now differentiate both sides with respect to 't'

OpenStudy (azureilai):

So would that be \[\frac{ ds }{ dt }=2\pi x\]

OpenStudy (baru):

u have missed out something on RHS. i said differentiate with respect to 't'. you have to assume x is a function of t (x=f(t) )

OpenStudy (azureilai):

ah so \[\frac{ ds }{ dt }=2\pi \frac{ dx }{ dt }\]

OpenStudy (baru):

close...but still wrong. you have to use the 'chain rule' \[\frac{ds}{dt}=2\pi x \frac{dx}{dt}\]

OpenStudy (azureilai):

Ok. I see where I messed up. So in the next step, do I rearrange it and plug in numbers?

OpenStudy (baru):

yep, be careful, the question says surface area "decreases" so you have to put a minus sign :)

OpenStudy (azureilai):

Ok. Thank you very much